OpenStudy (anonymous):

Use the angle sum identity to find the exact value of each 1. Cos 105 degrees 2. Sin 195 degrees

OpenStudy (mathmate):

hints: 105 degrees = 60+45 degrees 195 degrees = 135+60 degrees. And use sin(x)=sin(180-x). For values of sin(45), cos(45), sin(60), cos(60), use the unit circle. The sum formula is cos(a+b)=cos(a)cos(b)-sin(a)sin(b) sin(a+b)=sin(a)cos(b)+cos(a)sin(b).